OpenStudy (anonymous):

The graph of a system of parallel lines will intersect at exactly 1 point. Always Sometimes Never

terenzreignz (terenzreignz):

Parallel lines... @allie_bear22 It's in their very definition :D

OpenStudy (anonymous):

"parallel" train tracks.. do they ever intersect?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so it would be never

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

there you go

terenzreignz (terenzreignz):

Yup

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thank you can you help me with another?

terenzreignz (terenzreignz):

aww... what the heck... Hit me :D

OpenStudy (anonymous):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 2x + y = 8 -x - 3y = -12

terenzreignz (terenzreignz):

Well, isolate x in the second equation, please... What isolate means~ manipulate the equation so that x stands alone on one side of the =sign

OpenStudy (anonymous):

okay wat next

terenzreignz (terenzreignz):

Isolate x first, and I'll tell you what's next :)

OpenStudy (anonymous):

im not sure how to isolate it

terenzreignz (terenzreignz):

There's a little technique in equations called "transposition" ever heard of it?

OpenStudy (anonymous):

yea

terenzreignz (terenzreignz):

that involves taking a term, and bringing it to the other side, but changing its sign in the process.

OpenStudy (anonymous):

am i supposed to do it to both of the equations or just 1?

terenzreignz (terenzreignz):

Just the second equation. Do it until you can get x to stand alone (by alone, I also mean no -sign next to it)

OpenStudy (anonymous):

3y-12=-x

terenzreignz (terenzreignz):

Yeah, good... but there is still a -sign next to the x.

OpenStudy (anonymous):

because its negitive

terenzreignz (terenzreignz):

Yeah, so what you can do in this case, you can multiply both sides of the equation by -1.

OpenStudy (anonymous):

-3y-12=x

terenzreignz (terenzreignz):

You forgot the -12... why is it still negative? :P

OpenStudy (anonymous):

-3y+12=x

terenzreignz (terenzreignz):

There. so you now have x = 12 - 3y Now substitute this value for x in the x in the first equation, simplify, and that's your answer.

OpenStudy (anonymous):

2=12-3y

terenzreignz (terenzreignz):

2x + y = 8 Could you do that again? Replace x with (12 - 3y)

OpenStudy (anonymous):

(12-3)=2?

terenzreignz (terenzreignz):

\[\huge 2\color{red}x + y = 8\]\[\huge \color{red}{x = 12 - 3y}\]

terenzreignz (terenzreignz):

Replace.

OpenStudy (anonymous):

(12-3)+y=8 or 12-3y=x

terenzreignz (terenzreignz):

\[\huge 12-3\boxed{\color{red}y}\]Do not leave out these things...

OpenStudy (anonymous):

okay wat next

terenzreignz (terenzreignz):

Not yet, not until you do this correctly. 2x + y = 8 x = 12 - 3y Now replace x in the first equation.

OpenStudy (anonymous):

12-3y

terenzreignz (terenzreignz):

Yes... replace x in the first equation with that.

OpenStudy (anonymous):

12-3y=8

terenzreignz (terenzreignz):

Nope.... I said replace x, not replace the entire left side...

OpenStudy (anonymous):

2(12-3) + y = 8

terenzreignz (terenzreignz):

I told you not to leave this out... \[\huge 12-3\boxed{\color{red}y}\]

OpenStudy (anonymous):

2(12-3y)=8

terenzreignz (terenzreignz):

It was 2x + y = 8 not 2x = 8

OpenStudy (anonymous):

2x+(12-3y)=8

terenzreignz (terenzreignz):

No. 2x + y = 8 replace x with 12 - 3y (I'm really not proceeding from here until you get this step right, this is the most important step)